EDITORIAL

Terry Goodman

Welcome to a new year of MJMS. As we begin the eighth volume of MJMS, you will notice that the format of our journal has changed significantly. First MJMS is now a little different in size and has a new cover design. We have also chosen to use a better, “slick” paper for the cover. I want to thank Hang Chen and Curtis Cooper of CMSU for their excellent work in developing the new cover design. We believe these changes will make the journal more durable and readable.

We are considering the possibility for the future of adding a fourth issue each year. We would appreciate your input and feedback concerning such a change as well as any other suggestions you have concerning MJMS.

We now have an MJMS home page on the WEB where you can find information about the journal as well as past articles and problems. Our WEB address is:

http://153.91.1.102/˜ hchen/mjms/mjms.html

We hope you will find our home page useful and interesting.

You may recall that in my editorial of last fall I raised the issue of identifying the three greatest mathematicians of all time. We received several responses concerning this question and, as we expected, there was not complete agreement among the respondents. In the History of Mathematics course that Bob Kennedy and I taught, we identified Archimedes, Newton, and Gauss as the three greatest mathematicians. Obviously, there is no definitive answer to this question, but it is interesting to discuss the rationale for making such selections. For a very thoughtful and thorough discussion of this issue I would recommend “The Mathematician of the Millennium” by Richard L. Francis of Southeast Missouri State University. You can find this article in the Fall, 1991 issue of Consortium, the newsletter of the Consortium for Mathematics and Its Applications. We will continue to welcome input from our readers concerning this question. æ